# Dynamics of inviscid annular supersonic jets

### v1

@@ 4.1.2 Solving the Model
It is important to note that this approximation leads to an overestimation of $\mathcal{M}_{i}$ since, in real flow conditions, the Mach disk always forms downstream of the separation point.
Finally, the thrust of the control volume, $T_{w}$, can be determined by integrating the wall pressure along the nozzle from the throat to the separation point $x_{sep}$:
@@
The model is now complete and only requires the nozzle wall radius and Mach number profiles, which can be computed using methods like the Method of Characteristics (MOC). An example using nozzle characteristics from Jaunet et al. (2017) is presented in the following section. The acoustic Mach numbers of the vortex sheets and the inner radius satisfying the proposed model are depicted in Figure 16 against the fully expanded Mach number of the flow $M_{j}$. $M_{j}$ can be easily computed from the nozzle NPR using isentropic relations. The separation criterion from Stark (2005), with slight modifications to better align with the experimental observations of Jaunet et al. (2017), is employed. The decrease in $\mathcal{M}_{i}$ as the normal shock wave strengthens with downstream displacement of the separation, is in accordance with expectations. Conversely, neither $M_{a}$ nor $D_{i}$ display a monotonic behavior. This can be attributed to the nozzle wall profile approaching a near-horizontal orientation near the lip, resulting in saturation of thrust $T_{w}$. Another noteworthy behavior is the sudden decrease in $D_{i}$ as the separation approaches the nozzle lip, consistent with the behavior of overexpanded jets from TIC nozzles—akin to an approaching adaptation where the Mach disk diminishes or disappears.
@@
In conclusion, we have formulated a comprehensive model that derives essential parameters for feeding a simplified linear dynamical model of jet flows exiting from over-expanded nozzles, based on nozzle pressure and Mach number profiles. In the subsequent sections, we present dynamics obtained using this model for various flow configurations.
@@ 4.2 Stability properties of the flow with varying NPR
Using the previous simple model we obtained all the necessary parameters to build simplified base flows for $M_{j}$ varying in between $1.7$ and $2.2$, corresponding to expansion where resonances were observed. The aforementioned base flows are presented in figure 17 where the reader can see that the position of the inner mixing layer, as well as both the inner and annular flow velocity, vary with the chosen jet Mach number, their trends being in agreement with the results of figure 16. Regarding the thickness of the mixing layers, we used that obtained from the numerical results shown in figure 2 for all cases. This is obviously a strong simplification, but it is believed that given the relative narrow range of jet Mach number considered, the exact mixing layer shape may not drastically change.
@@ 4.2.1 K-H modes
The dynamical characteristics of the unstable K-H waves with respect to the base flows presented in the former paragraph are plotted in figure 18 for various frequencies. As expected from the previous results, both KH waves are unstable, the outter one being the most unstable for all flow regime and frequencies considered. The evolution of the growth rate of the inner KH wave seems rather in dependant of the jet Mach number until $M_{j}=2.0$. Above this expansion ratio, it is interesting to notice that the inner KH wave is more and more unstable, for all the frequencies considered. It is pertinent to recall that experimental observations have reported that resonances in TIC nozzles only manifest around $M_{j}=2.09$ (Jaunet et al. (2017)). Although it is possible that the proximity of the experimental and this theoretical values is coincidental, it is worth mentioning. As it might be that resonance in TIC nozzle may only be allowed when the inner mixer layer becomes sufficiently unstable.
Fig. 18: Growth rate of both the innner ( ) and outter ( ) KH modes for varying jet Mach number $M_{j}$ and Strouhal number.
We report in figure 19 the pressure eigenfunctions of both KH modes. The outter KH well localized at the maximum shear of the outter mixing layer and decays exponentially in the radial direction. For all the cases considered here, the inner KH-mode spatial support is more widely spread across the jet and shows a slower radial decay that the outter one. Overall, the KH waves computed in those cases show the same characteristics than previously. The reader must notice, however, that the spatial support of the KH waves computed does not depend on the Strouhal number nor the jet Mach number.
Fig. 19: Eigenfunction of both the innner ( ) and outter ( ) KH modes for varying jet Mach number $Mj$ and Strouhal number.
@@ 4.2.2 Guided jet modes
As far as guided jet modes is concerned, the expansion ratio of the nozzle plays a significant role in their dispersion relation, as can be seen in figure 20. At low $M_{j}$, only upstream propagating guided jet modes seems supported by the flow field. As depicted earlier in the study, this is and effect of the thinning of the annular jet providing the modes a behavior encountered in subsonic jetJordan et al. (2018); Towne et al. (2017). As soon as the overall speed of the increases, the GJM dispersion relation deforms towards what is expected in supersonic cases where the dispersion relation is constituted by downstream- and upstream-propagating branches. In any case, the results show that for all expansion ratio, the flow supports an upstream-propagating wave making resonance loops possible.
Fig. 20: Dispersion relation of the guided jet modes for varying nozzle expansion ratio.
@@ 4.2.3 Discussion
One may notice that for $M_{j}=2.09$, the dynamical model does not show any upstream propagating waves at the $St=0.2$, where a resonance was experimentally observed Jaunet et al. (2017); Bakulu et al. (2021). This means that it wouldn’t be possible with the current base flow and finite thickness models to predict a resonance at the correct Mach and Strouhal numbers in over-expanded flows, with a similar mechanism as depicted for screech like resonance. We believe that this has to be attributed to the numerous hypothesis and simplification that are embedded in these models. Nevertheless, the analysis conducted in this study revealed that the topology of the flow exiting a convergent divergent nozzle do support the necessary instability waves to produce a resonance (i.e. a downstream KH waves and an upstream GJM). The sensitivity of the waves dispersion relation and stability characteristics to the base flow, shown in the previous section, indicates that a more precise definition of the base flow would be necessary if a precise resonance prediction is needed. The reader may be referred to the work of Chow & Chang (1975) or Li & Ben-Dor (1998) for this purpose.

### v2

@@ 2.1.3 Solving the base flow model
It is important to note that this approximation leads to an overestimation of $\mathcal{M}_{i}$ since, in real flow conditions, the Mach disk always forms downstream of the separation point, hence at higher Mach numbers.
Finally, the thrust of the control volume, $F_{w}$, can be determined by integrating the wall pressure along the nozzle from the throat to the separation point $x_{sep}$:
@@
The model is now complete and only requires the nozzle wall radius and Mach number profiles, which can be computed using methods like the Method of Characteristics (MOC). An example using nozzle characteristics from Jaunet et al. (2017) is presented in the following section. The acoustic Mach numbers of the vortex sheets and the inner radius satisfying the proposed model are depicted in Figure 4 against the fully expanded Mach number of the flow $M_{j}$. $M_{j}$ can be easily computed from the nozzle NPR using isentropic relations. The separation criterion from Stark (2005), with slight modifications to better align with the experimental observations of Jaunet et al. (2017)) for this nozzle geometry, is employed. The decrease in $\mathcal{M}_{i}$ as the normal shock wave strengthens with downstream displacement of the separation, is in accordance with expectations. Conversely, neither $M_{a}$ nor $D_{i}$ display a monotonic behavior. This can be attributed to the nozzle wall profile approaching a near-horizontal orientation near the lip, resulting in saturation of thrust $F_{w}$. Another noteworthy behavior is the sudden decrease in $D_{i}$ as the separation approaches the nozzle lip, consistent with the behavior of overexpanded jets from TIC nozzles—akin to an approaching adaptation where the Mach disk diminishes or disappears.
@@
In conclusion, we have formulated a comprehensive model that derives essential parameters for feeding a simplified linear dynamical model of jet flows exiting from over-expanded nozzles. This model requires the a priori knowledge of the static pressure and Mach number profiles along the nozzle and a separation criterion. From this, all the necessary quantities to build a simplified mean annular supersonic mean flow for various nozzle regimes can be retrieved.
2.2 Finite thickness dynamical model
We propose to use the same approach as Mancinelli et al. (2020), modelling the jet using locally-parallel linear stability theory. All variables are normalised by the fully expanded jet diameter $D_{j}$, the ambient density and speed of sound $\rho_{\infty}$ and $c_{\infty}$, respectively. The Reynolds decomposition,
$q^{\prime}(x,r,\theta,t)=\overline{q}(r)+q(x,r,\theta,t),$ (13)
$q(x,r,\theta,t)=\hat{q}(r)e^{i(kx+m\theta-\omega t)}$ (14)
where $k$ is the streamwise wavenumber and $m$ is the azimuthal Fourier wavenumber. The non-dimensional pulsation $\omega=2\pi StM_{a}$ is computed via the acoustic Mach number of the jet $M_{a}=U_{j}/c_{\infty}$ and the Strouhal number of the flow $St=f{D_{j}}/{U_{j}}$, where $U_{j}$ and $D_{j}$ are the equivalent fully expanded velocity of the flow and diameter of the jet, respectively.
where $\gamma$ is the specific heat ratio for a perfect gas. The solution of the linear stability problem is obtained specifying a real or complex frequency $\omega$ and solving the resulting augmented eigenvalue problem $k=k(\omega)$, with $\hat{p}(r)$ the associated pressure eigenfunction. The eigenvalue problem is solved numerically by discretizing 15 in the radial direction using Chebyshev polynomials. A mapping function is used to non-uniformly distribute the grid points such that they are dense in the region of shear (Trefethen, 2000).
Although the model described above supports non-isobaric regime (Mancinelli et al., 2023), only isobaric jets will be studied in the following. Moreover, only the first non-axisymmetric fluctuating mode ($m=1$) is considered, for this azimuthal mode is the only one responsible for side-loads in overexpanded nozzle flows (Dumnov, 1996), and for resonances observed in experiments (Jaunet et al., 2017).
The flow dynamics is also studied without taking the nozzle walls into account. This corresponds to a free jet configuration and we are looking for pressure waves vanishing at infinity. This choice is motivated by the fact that, in our previous observations, the downstream- and uptream-propagating waves signatures were observed as being dominant far downstream of the nozzle, where the influence of the nozzle wall is negligible.
2.3 Convergence of the linear stability calculation
We present in figure 5 the convergence of the eigenvalue spectrum with respect to the number of Chebychev collocation points $N$ and the Strouhal number. We indicated in figure 5 the position of the waves of interesrest in this paper, namely the guided jet modes (GJM) and the two unstable waves ($k_{i}<0$) that lie in the $k_{r}>0$ portion of the spectrum and are downstream travelling. These modes will be named inner and outer Kelvin-Helholtz (KH) waves in the document for they have spatial support with a maximum located on both the inner and the outer mixing layers, as expected for Kelvin-Helmholtz (KH) waves. This denomination seems logical although it will be shown later that their spatial supports shows some differences, especially regarding their decay in the radial direction. This might need a separate discussion by itself, but it is out of the scope of the curent study.
The GJM, on the other hand, lie in the $k_{r}<0$ region and on the real axis, they hence are neutral wave with negative phase speed. With varying frequency, the position of these modes changes. The GJM moves along the real axis and might separate from this axis and become evanescent. That is what wan be seen in figure 5 right, where two GJM can be seen near by the $(-5,\pm 1.5)$ position. For a more precise discussion on these modes, we refer the reader to Towne et al. (2017).
The overlay of the different symbols in figure 5 shows that the convergence is more difficult to reach at high frequencies, especially for the Kelvin-Helmholtz modes, but the spectrum seems to have converged above $N=301$ points for the frequencies of interest in this paper. Therefore, a total number of $N=401$ collocation points is used for all the computations in this study.
Fig. 5: Convergence of the eigenvalue spectrum with respect to the number of Chebychev collocation points for $St=0.1$ (left) and $St=0.4$ (right).
3 Dynamics of annular supersonic jets
Contrary to a single stream jet flow, the base flow of a annular jet requires more parameters to be perfectly described. The characteristic features of the base flow (mixing layer thickness or position) may evolve with the nozzle pressure ratio as well and, therefore, their relative influence on the stability spectrum needs to be evaluated before one may explore the conditions for resonance. For this purpose, we use the analytical profile described above and explore the influence of the main parameters on the stability properties of the jet.
3.1 Effect of radial position
The influence of the relative positions of the mixing layers is first studied by assuming a constant reference velocity levels in the subsonic core and the supersonic stream. We perform linear stability analyses varying the position of the inner mixing layer and keeping the one of the outer mixing layer constant, as presented in Figure 6. The inner, annular, and external Mach numbers are chosen to approximately match the numerical simulations downstream of the first Mach disk: $M_{i}=0.6$, $M_{a}=1.4$, and $M_{e}=0$. The density profile was computed using the Crocco-Busemann relation (Michalke, 1984). A total of 36 base flows is used to finely study the changes in the corresponding linear dynamics.
Typical eigenspectra obtained at $St=0.2$ and $m=1$ are presented in Figure 7. On the right-hand side of the figure ($k_{r}>0$), two groups of unstable eigenvalues (i.e. belonging to the $k_{i}<0$ half plane) can be observed. They correspond to the Kelvin-Helmholtz (K-H) modes of both the inner and outer mixing layers, as was found by Dahl & Morris (1997b) for supersonic IVP jets. The inner K-H mode is the least unstable of the two. The vertical ($k_{r}=0$) and horizontal ($k_{i}=0$) parts of the continuous acoustic branch are clearly visible, with a relatively weak influence of $R_{i}$ on the loci of the corresponding eigenvalues. Propagating and evanescent guided jet modes (Towne et al., 2017) can also be identified in the $k_{r}<0$ half plane. Thus, the annular supersonic jets appear to exhibit similar features to those of a classical top-hat supersonic jet. In the following, we will focus on extracting the effect of the inner mixing layer position on the characteristics of these waves.
3.1.1 Kelvin-Helmholtz mode
In Figure 8, we plot the growth rate evolution of both K-H modes as a function of $R_{i}$ for various Strouhal numbers. The outer K-H mode is always more unstable than the inner one for all the configurations investigated in this study. This is expected considering that the velocity gradient is stronger for the outer mixing layer than for the inner one. The outer mode becomes even more unstable as $R_{i}$ increases, although at high frequencies its growth rate shows a plateau at low values of $R_{i}$. On the other hand, the inner mixing layer instability shows an optimal growth rate at specific $R_{i}$ values. This trend suggests that there may exist specific base flows and frequencies for which the inner and outer K-H modes possess equivalent growth rates. In some cases, the inner mode can even become more unstable than the outer one. However, this is more likely to occur at higher frequencies and might not be relevant for the initial nozzle flow problem, where resonance at lower frequencies is observed.
The pressure eigenfunctions associated with the two identified K-H modes are presented in Figure 9 for $St=0.2$. As expected, the eigenfunction is maximum at the mixing layer location and is zero at the centerline, as we focus on the $m=1$ anti-symmetric wavenumber. The outer K-H mode does not seem to be significantly affected by the inner mixing layer, unlike the inner mode whose support spreads further to the outside the outer mixing layer as the two mixing layers get closer. Interestingly, the radial support of the inner K-H wave decays less rapidly than that of the outer one. This allows the inner mode to be detected quite far from the jet and enables it to exchange energy with feedback waves outside of the supersonic annular region. This aligns well with the observations of Bakulu et al. (2021), who found that the inner mixing layer supports most of the perturbations associated with the resonance process in their flow.
3.1.2 Guided jet modes
Guided jet modes have been shown to play a significant role in the resonance of screeching and impinging jets (see, for example, Gojon et al. (2018); Jaunet et al. (2019); Mancinelli et al. (2021); Edgington-Mitchell et al. (2018)). Their dispersion relation is a key characteristic that explains cut-on and cut-off frequencies, the transition from axisymmetric to helical modes, and improves the model predictions (Mancinelli et al., 2019, 2021; Nogueira et al., 2022). As explained previously, their eigenvalues lie on the real axis of the spectrum until they become evanescent. Figure 7 shows that the annular jet configuration, as encountered in the exhaust of over-expanded CD nozzles, also supports such type of waves. In the $(k_{r},St)$ plane, this occurs at the summit of a given branch. At supersonic speeds, jets support both upstream propagating, denoted $k_{p}^{-}$, and downstream propagating, denoted $k_{p}^{+}$, GJM. This distinction is made based on the sign of their group velocity $u_{g}=\frac{\partial\omega}{\partial k}$, hence corresponding to the local slope of their dispersion relation presented in figure 10. The $k_{p}^{-}$ can be seen close to the acoustic waves dispersion relation and is the mode that can carry energy in the upstream direction. The dispersion relation of these waves is plotted in Figure 10 in the $St-k_{r}$ plane for various positions of the inner mixing layer.
Fig. 10: Neutral mode dispersion relation for various positions of the inner mixing layer: $R_{i}=0.05$ (top left),$R_{i}=0.1$ (top right),$R_{i}=0.2$ (bottom left),$R_{i}=0.35$ (bottom right). The dash-dotted line correspond to the acoustic waves dispersion relation.
As expected, at each $R_{i}$, the neutral waves form families of modes hierarchically ordered by their radial supports: higher frequencies correspond to higher radial orders (i.e., the number of nodes and antinodes, see Tam & Hu (1989)). As seen, the position of the inner mixing layer has a strong impact on the neutral modes’ dispersion relation: as $R_{i}$ increases, the frequencies at which the modes are encountered also increase. Furthermore, the group velocity of the $k_{p}^{+}$ mode decreases with increasing $R_{i}$. This can be understood by recognizing that for $R_{i}=0$ or $R_{i}=R_{e}$, the base flow is close to a supersonic top-hat jet or a subsonic one, respectively. As can be seen, both group and phase velocities of the guided jet modes decrease with increasing $R_{i}$. This is expected from the decrease of the overall average speed of the jet when $R_{i}$ increases. The neutral modes’ behavior seems to lie between what is expected for a subsonic $M=M_{i}<1.0$ jet and a supersonic $M=M_{a}>1.0$ one.
The eigenfunctions associated with the upstream propagating neutral modes are presented in Figure 11 for $St=0.1$ and $St=0.2$. As expected, the radial supports of these modes share common features with Bessel functions. At low Strouhal numbers, the eigenfunctions show one antinode, while for higher frequencies, more antinodes can be observed. Surprisingly, very little change in the radial support is observed for varying $R_{i}$. This is due to the fact that these modes are duct-like modes, so that their support is mostly independent of the base flow, in contrast to their wavenumber. It is important to notice that they also have support outside the jet, similar to the inner K-H wave. This allows these waves to interact and for the upstream mode to carry energy upstream, potentially closing a feedback loop.
3.2 Effect of mixing layer thickness
In order to explore the influence of the thicknesses of the mixing layers on the stability properties of the flow, we vary the base flow as presented in Figure 12. Both mixing layers are centered around $D_{j}/2$ and $D_{j}/4$. The jet bulk velocities are chosen to be constant and comparable to those observed in the nozzle after the first Mach disk for the case previously studied at $M_{j}=2.1$.
Fig. 12: Base flows used to study the effect of mixing layer thickness.
3.2.1 Kelvin-Helmholtz mode
Figure 13 shows the growth rate of both KH modes with varying thicknesses and Strouhal numbers. As expected, larger mixing layer thicknesses result in lower growth rates of the KH mode. For the base flows and frequencies considered here, the inner mixing layer is more stable than the outer one. This is due to their difference in velocity gradient. The outer mixing layer separates a supersonic stream from a quiescent atmosphere, while the inner one separates two rapid flows. Interestingly, for relatively high frequencies (above $St=0.2$), increasing the thickness stabilizes the outer mixing layer more quickly than the inner one. Thus, there might be flow configurations for wich both KH modes exhibit comparable growth rates. Depending on the frequency and base flow, the dynamics may not be entirely dominated by the outer KH mode, as one might initially assume.
The associated spatial profiles of the KH modes mentioned above are presented in Figure 14 for different Strouhal numbers and varying mixing layer thicknesses. We observe similar results to the previous study: for the thinnest mixing layer, the inner and outer KH modes exhibit maximum pressure fluctuations at the locations where the base flow exhibits the maximum levels of velocity gradients. For all frequencies, thickening the mixing layer tends to spread the spatial support of the inner modes towards the outer mixing layer by increasing the pressure amplitude near the jet boundary. At the highest frequencies studied here, the signature of the outer KH modes also spreads onto the inner mixing layer. As indicated earlier by their relative growth rates, the entanglement of spatial support of the KH modes with increasing mixing layer thickness might make it difficult to distinguish their relative signatures in the fluctuating pressure field. Nonetheless, as mentioned in the previous section, the inner mixing layer KH mode appears to have a slower radial decay compared to the outer one. This supports the findings of Bakulu et al. (2021), which show that at the resonance frequency, the downstream energy-carrying mode was supported by the inner mixing layer. The wide radial extent of the eigenfunction, as shown here, also suggests that the inner KH mode can exchange energy with guided-jet modes or acoustic modes, whose support can be located outside the jet.
3.2.2 Guided jet modes
The dispersion relation of the guided jet modes for varying mixing layer thicknesses is now presented in figure 15. Unlike the dispersion relation of K-H modes, the dispersion relation of GJM is only little affected by thicknesses of mixing layers. There is a slight variation in the domain of existence of the upstream-traveling modes between the branch and saddle points, as pointed out by Towne et al. (2017). The primary effect of thickening the mixing layers is on the group velocity ($\partial\omega/\partial k$), with thicker mixing layers leading to lower group velocities, as one would expect.
@@ 3.3 Stability properties of the flow with varying NPR
Using the simple base flow model described in section 2.1.2, we obtained all the necessary parameters to build simplified base flows for $M_{j}$ varying in between $1.7$ and $2.2$, corresponding to expansion regimes where resonances were observed for the TIC nozzle geometry previously studied (Jaunet et al., 2017). The aforementioned base flows are presented in figure 17 where the reader can see that the position of the inner mixing layer, as well as both the inner and annular flow velocity, vary with the chosen jet Mach number, their trends being in agreement with the results of figure 4. Regarding the thickness of the mixing layers, we used that obtained from the numerical results shown in figure 2 for all cases. In spite of this simplification, it is believed that given the relative narrow range of jet Mach number considered, the exact mixing layer shape may not drastically change.
@@ 3.3.1 K-H modes
The dynamical characteristics of the unstable K-H waves with respect to the base flows presented in the former paragraph are plotted in figure 18 for various frequencies. As expected from the previous results, both KH waves are unstable, the outer one being the most unstable for all flow regime and frequencies considered. The evolution of the growth rate of the inner KH wave seems rather independent of the jet Mach number until $M_{j}=2.0$. Above this expansion ratio, it is interesting to notice that the inner KH wave is more and more unstable, for all the frequencies considered. It is worth noting that this value is very close to the value of $M_{j}=2.09$ around which resonances were experimentally observed (Jaunet et al., 2017). The proximity of these experimental and theoretical values is very striking and still supports the idea that resonance in TIC nozzle may only be allowed when the inner mixing layer becomes sufficiently unstable.
Fig. 18: Growth rate of both the innner ( ) and outer ( ) KH modes for varying jet Mach number $M_{j}$ and Strouhal number.
We report in figure 19 the pressure eigenfunctions of both KH modes. The outer KH well localized at the location of maximum shear in the outer mixing layer and decays exponentially in the radial direction. For all the cases considered here, the inner KH-mode spatial support is more widely spread across the jet and shows a slower radial decay that the outer one. Overall, the KH waves computed in those cases show the same characteristics than previously. The reader must notice, however, that the spatial support of the KH waves computed does not depend on the Strouhal number nor the jet Mach number.
Fig. 19: Eigenfunction of both the innner ( ) and outer ( ) KH modes for varying jet Mach number $Mj$ and Strouhal number.
@@ 3.3.2 Guided jet modes
As far as guided jet modes is concerned, the expansion ratio of the nozzle plays a significant role in their dispersion relation, as can be seen in figure 20. At low $M_{j}$, the slope of the dispersion relation is negative and do not show any extrema, hence only upstream propagating guided jet modes seems supported by the flow field. As depicted earlier in the study, this is an effect of the thinning of the annular jet providing the modes a behavior encountered in subsonic jet (Jordan et al., 2018; Towne et al., 2017). As soon as the overall speed of the jet increases, the GJM dispersion relation deforms towards what is expected in supersonic cases where the dispersion relation is constituted by both downstream- ($\frac{\partial\omega}{\partial k}>0$) and upstream-propagating ($\frac{\partial\omega}{\partial k}<0$) branches. In any case, the results show that for all expansion ratio, the flow supports an upstream-propagating wave making resonance loops possible.
Fig. 20: Dispersion relation of the GJM for varying nozzle expansion ratio. The circles represents the dispertion relation for $M_{j}=2.09$.
@@ 3.3.3 Discussion
According to figure 20, the dynamical model does not show any upstream propagating waves for $M_{j}=2.09$ at $St=0.2$, although a resonance was experimentally and numerically observed (Jaunet et al., 2017; Bakulu et al., 2021). This means that it wouldn’t be possible with the current base flow and finite thickness models to predict a resonance at the correct Mach and Strouhal numbers in over-expanded flows, with a similar mechanism as depicted for screech like resonance. This is mainly due the numerous hypothesis and simplification that are embedded in these models.
The sensitivity of the waves dispersion relation and stability characteristics to the base flow, shown in the previous section, indicates that a more precise definition of the base flow would be necessary if a precise resonance prediction is needed. The reader may be referred to the work of Chow & Chang (1975) or Li & Ben-Dor (1998) for this purpose.
The use of a parallel, isobaric flow model might also be cause of this inadequation between the model and the observations. Indeed, the flow exiting the nozzle is not parallel. For example, as for underexpanded jets, the shock cell network modulates the mean flow in the axial direction aand it has been shown to be an important feature in screeching jets (Nogueira et al., 2022).
Despite these limitations, the analysis conducted in this study revealed that the topology of the flow exiting a convergent divergent nozzle do support the necessary instability waves to produce a resonance (i.e. downstream KH waves and upstream GJM).